### v1

# FedCLEAN: byzantine defense by CLustering Errors of Activation maps in Non-IID federated learning environments

Abstract
Federated Learning (FL) enables clients to collaboratively train a global model using their local datasets while reinforcing data privacy. However, FL is susceptible to poisoning attacks. Existing defense mechanisms assume that clients’ data are independent and identically distributed (IID), making them ineffective in real-world applications where data are non-IID. This paper presents FedCLEAN, the first defense capable of filtering attackers’ model updates in a non-IID FL environment. The originality of FedCLEAN is twofold. First, it relies on a client confidence score derived from the reconstruction errors of each client’s model activation maps for a given trigger set, with reconstruction errors obtained by means of a Conditional Variational Autoencoder trained according to a novel server-side strategy. Second, we propose an ad-hoc trust propagation algorithm based on client scores, which allows building a cluster of benign clients while flagging potential attackers. Experimental results on the datasets MNIST and FashionMNIST demonstrate the robustness of FedCLEAN against Byzantine attackers in non-IID scenarios and a close-to-zero benign client misclassification rate, even in the absence of an attack.
Keywords:
Poisoning attacks Byzantine attacks Non-IID Federated Learning Conditional variational autoencoders Clustering
1 Introduction
In recent years, Federated Learning (FL) has attracted considerable interest as a privacy-preserving machine learning paradigm [40, 32, 25] in several domains, such as healthcare [36], transportation [36], and personal devices [28]. In a typical FL scenario, several clients collectively contribute to training a global model. At first, a central server initiates the global model and sends it to the clients so that they can train it locally on their private data. Model updates are sent back to the server, which aggregates them into a new global model. These two steps are repeated until the global model converges. As clients’ data remain stored locally, privacy is ensured [42]. However, even though FL systems preserve privacy, their decentralized nature makes them vulnerable to poisoning attacks, where one or several malicious clients try to corrupt or prevent the training session by sending malicious updates [50]. Some defenses try to mitigate these attacks using secure aggregation algorithms with the idea of reducing the impact of bad updates [41, 8, 26]. Nevertheless, these solutions remain limited, being, for instance, not resistant to attacks that can destroy the global model after just one round of a successful attack, like Byzantine attacks [12]. This has led to the proposal of FL defense systems devoted to the detection of malicious updates before the aggregation stage. Solutions proposed to date [24, 11, 10, 46, 6] have been proven to be Byzantine attack-robust under the hypothesis of Independent and Identically Distributed (IID) data, i.e., when training data is homogeneously distributed among clients in terms of the number of samples and class balance. However, such systems do not perform well in non-IID contexts [2], which are more likely encountered in real-life scenarios. In the following, we introduce these solutions and their limitations, then present our contribution: FedCLEAN, the first FL defense system adapted to non-IID contexts.
1.1 Related Works
As stated above, the first defenses against malicious updates were the development of robust aggregation algorithms. FedAvg [41] is one of the earliest and most commonly used aggregation methods in FL. It consists of using a weighted average of the updates sent by clients to actualize the global model. Clients’ dataset sizes are used as weights. Doing so dilutes the effect of sparse attackers but is not sufficient when their number increases. It is also not robust to FL attacks that amplify updates to dominate the global average [48], and ineffective in non-IID scenarios where an attacker has a larger dataset compared to benign clients. GeoMed [26] uses the geometric median of clients’ updates as a global update. KRUM [8] uses one of the local updates to generate a global model update. The chosen local update minimizes the sum of distances to its nearest neighbors. Other aggregation schemes have been proposed since [38, 34, 1, 23], but they focus on training performance rather than security. All of these methods have been shown to be ineffective under Byzantine attacks [24].
To better respond to attacks, "detect and remove" defense systems have been recently proposed. They aim at excluding malicious updates before the aggregation step. One first class of methods regroups systems which rely on anomaly detection techniques based on autoencoders. Li et al. [37] were the first to propose such a defense. It trains a copy of the global model on the test dataset on the server side, while training at the same time a variational autoencoder (VAE) on surrogate vectors randomly sampled from the weights of this model at each epoch. Although this solution performs better than aggregation-based defenses under Byzantine attacks, it requires training a VAE on an entire server-side dataset to achieve good results. This is quite unrealistic in real-life scenarios, as the server most likely won’t have access to a significant training dataset. FedCVAE [24] later proposed an improvement using an untrained conditional variational autoencoder (CVAE). When the central server receives updates from the clients, the defense system computes the geometric median (GM) of surrogate vectors from each update. Then it normalizes these vectors by subtracting their GM value. These normalized vectors are then inputted into the CVAE, conditioned on the current round number in the FL training. As the CVAE is not pre-trained, only inputs near zero yield reconstruction errors close to zero, and vice versa. FedCVAE shows some defense results in a non-IID scenario but only against attacks exaggerated in magnitude and thus fairly easy to detect. More recently, FedCAM [6], a CVAE anomaly detection-based approach, achieves the latest state-of-the-art results against Byzantine attacks in IID scenarios. Rather than using model weights, it works with their activation maps for a given trigger set on which a CVAE is trained with GeoMed as a regularization. However, FedCAM only works in IID scenarios, as we will see in Section 5.
A third class of methods, as introduced by FedGuard [11], aims at exploiting clients’ training data for detection without disclosing it. FedGuard [11] asks each client to train a local CVAE on its own dataset in addition to the main task and to send to the server both its model updates and the decoder part of its CVAE. The server then uses all client decoders to generate sample data to evaluate each received local model update’s performance. Then for the aggregation step, it only keeps models that achieve a certain accuracy. This solves the reliance on an auxiliary dataset for CVAE training. However, it incurs significant computing overhead and demands heavy client resources. In addition, this approach also depends on the clients having enough data to train a generative model. This is achievable in an IID scenario with a small number of clients but unlikely in a non-IID scenario where datasets vary in size and content. More importantly, this solution opens up clients to data leakage from the generative decoders through, for instance, membership inference attacks [49].
A last class corresponds to clustering-based approaches. They have shown promising results on anomaly detection without needing autoencoders. Among recent ones, we can cite DeepSight [46]. It uses a model inspection technique and a clustering scheme to create clusters of updates, and to pick out those with the best scores. DeepSight successfully stops backdoor attackers [27], even in the non-IID case, by selecting groups with minimal attacker influence. However, it does not work for Byzantine attacks where a single successful breach is sufficient to make the global model convergence fail. More recently, FLEDGE [10] used a combination of cosine similarity and clustering to filter out attacks. It provides results against Byzantine attacks but only a fraction of their data is non-IID.
1.2 Contribution
In this work, we propose FedCLEAN, an extension of FedCAM, a solution introduced by Bellafqira et al. [6]. Similar to FedCAM, FedCLEAN analyzes the GeoMed normalized CVAE reconstruction errors of clients’ activation maps for given inputs of a trigger set. Clients with high reconstruction errors are rejected based on an anomaly detection strategy. While FedCAM achieves top byzantine robustness compared to concurrent methods in IID scenarios [6], it fails to achieve good results in the context of non-IID data, as we will see in Section 5. We identified two main issues causing this failure, which we address with FedCLEAN. First, the CVAE training method proposed in FedCAM has poor generalization in non-IID contexts, as client updates are more heterogeneous. Specifically, clients with the most unique data often fall out of distribution for CVAE and end up with similar error scores to attackers. To overcome this issue, FedCLEAN employs a different CVAE training process that accounts for the temporal evolution of FL training with warmup epochs and better captures the statistical properties of benign updates using an annealing technique and layer selection. This ensures that all benign client activation maps, even outliers, are correctly reconstructed. Second, FedCAM relies on the mean of reconstruction errors as a dynamic error threshold. This approach risks allowing the selection of some attackers’ updates for the aggregation step if the average is high enough. It is important to note that retaining even one Byzantine attacker in a round is sufficient to destroy the global model. Moreover, this strategy also impacts the convergence or accuracy of the global model in the absence of attackers, as it eliminates around half of the participants in each round. In the non-IID case, clients are less redundant compared to IID data. To address this problem, FedCLEAN introduces a clustering algorithm that iteratively aggregates and concentrates benign candidates into a trusted cluster, enabling a more refined selection process. The rest of this paper is organized as follows. Section 2 provides an overview of FL technical background, FL attacks, and useful components of our solution. Section 3 presents the federated context we considered and its threat model. Section 4 introduces our proposal, FedCLEAN, followed by Section 5, which details our benchmark and provides an analysis of the experimental results. Finally, we conclude the paper in Section 6.
2 Background
2.1 Federated Learning (FL)
2.1.1 A Common FL Scenario
In this work, we considered an FL setting where $N$ clients $\{C_{k}\}_{k=1..N}$ with their datasets $\{D_{k}\}_{k=1..N}$ participate in a training session orchestrated by a central server $S$ to train a global model $M_{G}$. During each FL round $t$, $S$ scatters the current global model $M_{G}^{t}$ to a subset of $U$ randomly selected clients. Each client updates the received model on its local dataset for a number of local epochs and sends back the updated model to $S$. The server then uses an aggregation algorithm such as FedAvg [41] to get an updated version of the global model $M_{G}^{t+1}$ from the clients’ updated models $\{M_{u}^{t}\}_{u=1..U}$. The above training steps are then repeated until convergence or a stopping criterion is met.
2.1.2 Non-IID Data in Federated Learning
One major statistical challenge in FL is the so-called Non-Independent and Identically-Distributed (Non-IID) hypothesis [33, 55]. Indeed, in most real-life applications [19], local datasets are client-specific, and data collection processes are often inconsistent across clients, resulting in data heterogeneity. In the literature, four main heterogeneity types are identified [53]:
  • Label skew: Where class labels are distributed differently among clients.
  • Feature skew: Where clients have differently-distributed features for the same class.
  • Quality skew: Where data quality is inconsistent across clients.
  • Quantity skew: Where the amount of local data may be unbalanced across clients.
In practice, feature and quality skew are intrinsic properties of data and are hard to simulate for an experiment if the dataset does not provide a natural partition. So, most works that address Non-IID FL use a combination of label and quantity skew by choosing a distribution of data over clients that ensures heterogeneity.
2.1.3 Attacks in Federated Learning
Federated Learning is known to be vulnerable to several attacks [33]. During training, these attacks involve an adversary that compromises one or several clients and tries to use them to disrupt the global model. Herein, two main approaches can be distinguished: "model poisoning", where the attacker can directly manipulate the compromised clients’ models and send malicious updates to the server; and "data poisoning", where the attacker indirectly manipulates learning by corrupting the clients’ local datasets. Based on the attacker’s objective, these attacks can further be divided into two sub-classes:
  • Untargeted poisoning attacks - where the attacker’s goal is to drop the global model accuracy on any future inputs by preventing its convergence. It is generally very hard for a machine learning model to recover from a bad convergence, and one attack in an unprotected FL round is often sufficient for an untargeted poisoning attack to succeed. Such attacks are also referred to as "Byzantine attacks".
  • Targeted poisoning attacks - also called "Backdoor attacks", where the attacker aims at preserving the accuracy of the model on its main task while simultaneously training it on a second malicious task, introducing a backdoor. The most common form of backdoor attacks involves inserting a trigger in the training dataset and training the model to output a specific response when the trigger data are given as an input.
It is important to notice that making a successful backdoor attack is very hard in Non-IID scenarios [27]. Controlling just some of the FL clients with non-IID data, the attacker generally has access only to a subpart of the training data, which in some cases is not sufficiently representative for injecting a robust backdoor to benign clients’ updates [39]. Moreover, the backdoor updates are easily overwritten by benign clients’ ones, as shown in [39]. This is the reason why in this work, we decided to focus on defending against Byzantine attacks, which currently pose a greater threat in FL.
2.2 Conditional Variational Autoencoders (CVAE)
Figure 1: Architecture of a Conditional Variational Autoencoder (CVAE), where $x$ and $\hat{x}$ are the original and the reconstructed samples, respectively. $y$ is the condition, and $\mu$ and $\sigma$ are trainable parameters of the CVAE latent space.
Autoencoders (AEs) [30] are a class of unsupervised generative neural networks that aim to reconstruct a given input from a latent representation encoding. AEs are made of: i) an encoder $f$ that maps input vectors from the initial feature space to a lower-dimensionality representation space called the latent space; and ii) a decoder that retrieves input vectors from their latent representations. The objective of an autoencoder is to minimize the reconstruction error (RE) between the initial and reconstructed vectors. They have been historically used for dimensionality reduction [4, 30], model pre-training [20, 7], and generative tasks [52]. An important property of AEs is that, when applied to out-of-distribution vectors that deviate from the original training distribution, these deviations are amplified in the latent space, and much higher reconstruction errors are observed. This makes AEs considered as a valuable tool for anomaly detection in a wide range of applications [43].
Autoencoders having a fixed-dimensional latent space show some limitations for datasets with varying complexities. This has motivated the introduction of Variational Autoencoders (VAEs) [35] with a variational latent space based on Bayesian inference methods [18]. Conditional-VAEs (CVAEs) [45] are an extension of VAEs that leverages conditional probabilities to guide the encoding and decoding process following a provided condition. In practice, what differentiates a CVAE from a traditional VAE is the addition of a condition vector $y$ as an input of both the encoder and decoder (see Fig. 1). Thus, CVAEs enhance the capabilities of VAEs by making them compatible with supervised learning and allowing them to work in a task-guided manner [16]. They have been extensively used for a number of applications, including anomaly detection and detecting attackers in federated learning [24, 6, 11].
2.3 GeoMed and Activation Maps
In a Euclidean space, the geometric median (GeoMed) of a set of points is a point minimizing the $L_{2}$ distance to all the points in the set. GeoMed is a robust approximation of the centroid of a given set of points under the presence of outliers [26]. This has led to its extensive use in Byzantine-robust federated learning as a robust aggregation method [13] or to detect attackers [24].
In a typical deep learning model, the data is processed through feed-forward layers. Each hidden layer takes in the previous layer’s output, applies a linear transformation parametrized by a set of weights (e.g., a convolution or weighted sum), and then applies a nonlinearity function. The output of the layer’s weighted transformation is referred to as the activations of that layer. "Activation maps", or the set of activation maps extracted from a model for a given input, are characteristic of a model and the way it processes data. They have been used for applications such as explainability [47] and model watermarking [15, 5, 56]. It has been observed that in many real-life applications, the activation maps of a model tend to follow an intrinsic distribution [15, 44].
3 Threat Model and Defense Requirements
In this section, we present the threat model and security requirements considered in this work. In our FL scenario, we assume that the clients never share their datasets and only communicate their updated models to the server. Regarding the server, it has access to a sample of the testing set and the updated models received at each round (i.e., the server can inspect the inner layers and weights of the models clients send back after training).
3.1 Threat Model
3.1.1 Adversary’s Objective
In this work, we consider a "Byzantine" adversary $A$, who wishes to carry out Byzantine attack during the training phase of an FL model. To do so, $A$ controls a number of the FL clients who submit poisoned model updates to $S$ during the FL process with the objective of disturbing the convergence of the global model. Formally, $A$ attempts to fulfill the following objectives:
  • O1: If the attack succeeds, the global model’s accuracy on test data should collapse.
  • O2: The attacked model should not be able to recover from the accuracy drop by pursuing further training in the absence of the attacker.
3.1.2 Adversary’s Capabilities
As in [3, 46, 22, 11, 24, 10], we consider a strong adversary who has full control over $m<N/2$malicious" clients. $A$ has full access to these clients’ datasets, models, and communication channels. In particular, $A$ can conduct training using local datasets when clients are selected by the server, freely modify the weights of the resulting model updates, and send back modified updates to the server $S$ for aggregation. In accordance with Shannon’s Maxim, $A$ also has full knowledge of the aggregation method used by $S$ and of the defense this one deploys but cannot access or alter its parameters. They also have no access to other benign clients’ updates or datasets.
3.2 Security Requirements
To effectively defend against a Byzantine attacker, it is required that a filtering-based defense system $D$ satisfies the following requirements [46]:
  • R1: Block all malicious clients. As discussed above, the Byzantine adversary $A$ is able to break the model given one single successful round of attack. As a result, in order to be robust to such an adversary, $D$ must detect all malicious clients updates at any round.
  • R2: Do not disrupt training. $D$ should not negatively impact the training. In other words, $D$ must have high precision and not misclassify benign clients. It must also preserve the test accuracy achieved in the no-attack and no-defense scenario.
We will see that our proposal responds to all these objectives and security requirements.
4 Proposed defense: FedCLEAN
4.1 FedCLEAN Principle
The idea behind FedCLEAN is to analyze the normalized activation maps (NAMs), obtained by subtracting the GeoMed value from activation maps of one or several layers of $M_{G}$ for some given inputs referred as a trigger set $T_{s}=(X_{T},Y_{T})$, where $X_{T}$ and $Y_{T}$ are the set of samples and their labels, respectively. We assume that the NAMs obtained with malicious model updates will correspond to outliers that can be identified by means of an CVAE and GeoMed based anomaly detection strategy.
More clearly, before the FL process begins : FedCLEAN : i) makes a copy of the initial global model; ii) trains the copy locally for several epochs on a small dataset called the "trigger set"; iii) extracts the NAMs associated with the trigger set from different epochs of training; iv) trains a CVAE to reconstruct these NAMs, using their corresponding labels $Y_{T}$ as a condition. The training process is detailed in 4.2. By doing so, the CVAE learns the latent distribution of NAMs associated with each class.
Once trained, the CVAE is used by $S$ at each epoch to reconstruct for each client updates the NAMs corresponding to the trigger set. Malicious NAMs, which fall out of the learned distribution, result in high reconstruction errors. To discriminate malicious NAMs from benign ones, FedCLEAN makes use of a clustering procedure relying on an ad-hoc trust propagation algorithm. We discuss it in subsection 4.3.
To provide a formal description of FedCLEAN, let us consider a classification task and $N$ clients $\{C_{k}\}_{k=1..N}$. FedCLEAN operates according to the following steps :
  1. 1.
    FL and FedCLEAN Initialization
    1. (a)
      $S$ initializes the global FL model $M_{G}^{0}$ according to a common process. By assigning, for instance, random values to its weights.
    2. (b)
      Starting from $M_{G}^{0}$, $S$ trains its own version of the model: $M_{G}^{{}^{\prime}0}$, on its trigger set $T_{s}=(X_{T},Y_{T})$, where $X_{T}$ and $Y_{T}$ are the set of samples and their corresponding labels.
    3. (c)
      $S$ then trains the CVAE. To do so, it first extracts the activation maps $A_{S}$ of one or more layers from $M_{G}^{{}^{\prime}0}$ feeding it with the trigger set $T_{s}$. Then, it normalizes them using GeoMed such as: $\hat{A}_{S}=\sigma(A_{S}-\text{GeoMed})$ (where $\sigma$ is the sigmoid function). $S$ finally trains the CVAE on $\hat{A}_{S}$.
    4. (d)
      $S$ then sends $M_{G}^{0}$ to $U$ randomly selected clients ($U<N$), to initiate the first FL training round.
  2. 2.
    FL training process for each round $t>1$
    1. (a)
      Client training
      1. i.
        $S$ randomly selects $U$ clients $(U<K)$ to participate in the round $t$ and sends them the current global model $M_{G}^{t-1}$.
      2. ii.
        Each client $C_{u}$ trains $M_{G}^{t-1}$ on its local dataset $D_{u}$ for $n$ local epochs. The resulting updated model $M^{t}_{u}$ is next sent back to the server.
    2. (b)
      Secure FedCLEAN selection
      1. i.
        Inspection the activation maps for each client $u$: $S$ extracts the activation maps $A_{u}^{t}$ obtained by feeding the trigger set $T_{s}$ into $M^{t}_{u}$.
      2. ii.
        Scoring:
        • $S$ computes the activation maps’ GeoMed: $\text{GeoMed}(A_{1}^{t},...,A_{U}^{t})$ and normalizes the activation maps of each client: $\hat{A}_{u}^{t}=\sigma(A_{u}^{t}-\text{GeoMed})$.
        • $S$ feeds its CVAE with $\hat{A}_{u}^{t}$ conditioned on $Y_{t}$ and assigns to each client a reconstruction error score $\epsilon_{u}^{t}=\overline{MSE}(\hat{A}_{u}^{t},\tilde{A}_{u}^{t})$ where $\tilde{A}_{u}^{t}$ is the CVAE reconstruction of $\hat{A}_{u}^{t}$, and $MSE$ the mean square error function.
    3. (c)
      Filtering: Given the pairs $(M_{u}^{t},\epsilon_{u}^{t})$, $S$ performs our trust propagation clustering to filter out clients that are considered malicious and discards their updates.
    4. (d)
      FL Aggregation: Finally, $S$ aggregates updates from the remaining clients using FedAvg to generate $M_{G}^{t}$.
As exposed, the performance of FedCLEAN relies on two main components a CVAE and a trust propagation clustering algorithm used to discriminate malicious from benign updates. We come back to how they work in the following sections.
4.2 CVAE training constraints
As stated above, the CVAE model is trained by the server $S$ considering a simulated a benign model $M_{G}^{{}^{\prime}0}$ based on the server trigger set $T_{s}$. One of the constraints we assumed in building FedCLEAN is that $T_{s}$ is of very small size compared to the client data. To train the CVAE well, one thus needs to augment the training set. The solution we adopted is to use the activation maps (AMs) computed at different rounds of the training of $M_{G}^{{}^{\prime}0}$. But by doing so, we have to take into account the temporal drift of $M_{G}^{{}^{\prime}0}$ activation maps that impact CVAE generalization capabilities. In fact, AMs are more heavily affected by noise and very different at the very beginning of the training of $M_{G}^{{}^{\prime}0}$ than after a certain number of epochs. To overcome this issue, and as illustrated in Fig. 2 the CVAE training set is constituted of activation maps collected after some warmup epochs of the training of $M_{G}^{{}^{\prime}0}$. At each epoch, AMs are GeoMed normalized so as to better control AMs amplitude differences along the training of $M_{G}^{{}^{\prime}0}$. Indeed, with a GeoMed recomputed at each round, we have found that the distribution of the deviations from GeoMed remains stable. Anyway, even with a small trigger set, this technique allows to constitute a training set that captures the distribution of benign NAMs assuming certain hypotheses.
Figure 2: Construction of the CVAE training set.
A second issue we had to face when training the CVAE is that it is sensitive to "posterior collapse" [29, 9, 14], which causes it to converge to minima that ignore training inputs. More clearly, the traditional training of a CVAE, consists in maximizing the Evidence Lower Bound (ELBO) in the following loss:
$\mathcal{L}(x,y;\theta,\phi)=-\text{ELBO}(x,y;\theta,\phi)\\ =-(\mathbb{E}[\log p_{\theta}(x|z,y)]-\frac{1}{2}[\sigma_{\phi}(x)^{2}+\mu_{\phi}(x)^{2}-1-log(\sigma_{\phi}(x)^{2}))\\ =\text{MSE}(g_{\theta}(z,y),x)+D_{\text{KL}}(q_{\phi}(z|x)\|p(z))$ (1)
where: $g_{\theta}$ is the CVAE decoder; $q_{\phi}$ is the latent space distribution given by the CVAE encoder; $z$ is a latent representation sampled from this distribution; and, $(x,y)$ the CVAE input sample and its label; $D_{KL}$ is the Kullback-Leibler divergence measuring the divergence with a prior distribution $p$, in our case a standard normal distribution. As seen, this loss is constituted of a first term to minimize the CVAE reconstruction error and of a second to enforce a Gaussian structure for the latent space representation. In our context, NAMs give rise to numerically very small MSE values, so the optimizer focuses on minimizing the second term, leading to the so called "posterior collapse" issue. To solve this issue, it has been proposed to add a regularization weight to the divergence term in the loss function [29]. In our work, we adopted a KLD annealing approach with an adaptive regularization weight $\beta^{t}$ like [9]
$\mathcal{L}_{CVAE}=MSE+\beta^{t}D_{KL}$ (2)
where, $t$ is the current training epoch. CVAE starts to train with $\beta^{0}=0$ so that it learns to produce meaningful representations of the feature space and make predictions based on such representations. Then,$\beta$ is gradually increased by a fixed increase to enforce the structure of the latent space.
4.3 Clustering using trust propagation
A commonplace practice in CVAE anomaly detection and FL defenses [6, 24, 37] is to threshold error values to decide of an anomaly. Usually, this is the average reconstruction error of all clients. However because of this, in the absence of an attacker, nearly half of the clients even if they are all benigns are wrongly flagged as malicious. Well-adjusted static thresholds could allow for better filtering. In an IID scenario, such reference values could be implied from the error values of CVAE during training. However, in a Non-IID scenario, real client reconstruction error values cannot be deduced from the sample training scenario. We thus opted for a simple clustering algorithm instead.
To avoid systematically excluding benign clients, our clustering algorithm is able to automatically output either one or two clusters depending on whether malicious updates are present. This is a challenge as deciding the number of clusters based on data properties is an active problem in cluster analysis [21], and most algorithms require fixing the number of clusters beforehand as a user-defined parameter [21]. In addition, clustering algorithms vary widely in terms of "parameter reliance", i.e. the dependence on user-defined parameters and the tolerance to different ranges of such parameters. Our proposal has a single parameter that is interpretable (i.e. not rely on guessing).
It sequentially retrieves all benign clients participating in training, through what we call a "trust propagation" algorithm. It relies on a single relative parameter, which is a fraction in $[0,1]$ that can be selected in a data-agnostic manner, which it then uses to dynamically compute the distance measures it needs to perform clustering. Our idea relies on the fact that when reconstructing benign client activation maps, the measured reconstruction error is a residual from the randomness and uncertainty within CVAE. More clearly, when reconstructing out-of-distribution attacker NAMs, the reconstructed vector is an arbitrary response. Thus, the reconstruction error we measure is akin to measuring error to a random vector. As a result, assuming we are working with an autoencoder that has properly fitted benign NAMs distributions, we expect the error from the reconstruction of benign NAMs to be of a lower order of magnitude than those from anomalous NAMs. In other words: in the reconstruction error space, there is an identifiable relative error gap between the two.
Our algorithm works as follows. It sorts the client reconstruction errors, and selects the lowest reconstruction error client as the core of the benign client group. Then, in the reconstruction error space, if a client is within a certain distance $\delta$ of the last accepted client, he is added to the group. The iterations stop when the next client is at $d>\delta$ distance from the current client. We assume in that case that the gap between error values corresponds to that between attackers and benigns, and block all remaining clients. $\delta$ is computed dynamically as a proportion of the size of the error space for the current round, as $\delta=\lambda(max(\epsilon_{u}^{t})-min(\epsilon_{u}^{t}))_{0<u<U}$, where $\lambda\in[0,1]$ is a hyper-parameter fixed at the start of the defense system. The full pseudo-code for the trust propagation algorithm is detailed in Alg. 1.
Algorithm 1 Trust propagation
$\delta$, $clients\_sorted$, $sorted\_re$
$benign\_clients\leftarrow clients\_sorted[0]$
$i\leftarrow 0$
$i_{max}\leftarrow\text{len}(clients\_sorted)-1$
while ($i\neq i_{max}$) and ($sorted\_re[i+1]<sorted\_re[i]+\delta$do
     $benign\_clients.\text{append}(clients\_sorted[i+1])$
     $i\leftarrow i+1$
end while
return $benign\_clients$
5 Evaluation
5.1 Experimental setup
5.1.1 FL setting
We evaluate our proposition on Non-IID image classification tasks using two public datasets : MNIST [17] and FashionMNIST [51] datasets. The FL is orchestrated by a server that applies FedCLEAN as defense $D$ before each aggregation step with 100 total clients. 50% of clients are randomly selected to participate in training at each round, and we use FedAVG [41] to aggregate model updates. As global model $M_{G}$, we use a standard CNN architecture that achieves good results on both datasets in the absence of attackers. In all our experiments, we set 30% of the client population to be attackers.
5.1.2 Non-IID Data distribution
We distribute the samples of MNIST and FashionMNIST datasets among clients to synthesize a Non-IID data distribution. As mentioned above, we craft distributions that achieve label skew and quantity skew according two scenarios: like in [31] and [54] based on a latent Dirichlet distributions; and following our own custom Non-IID distribution where clients are only allowed to hold samples from two classes. More in details, these two secarios are such as:
  • Dirichlet distribution - we synthesize a population of $C$ clients by drawing a probability vector $\mathbf{q_{n}}\sim D_{C}(\alpha)$ for each class $n\in[1,N]$. Each client is allocated a portion $q_{n,c}$ of the total samples in class $n$. We set $\alpha=1$ [31].
  • Our custom distribution - we enforce the Non-IID condition originally proposed in [41] where each client only possesses samples from two classes. This is achieved by sampling the full datasets using an inverse law with a scale factor of $\alpha$ and an offset of $\gamma$. For the $c^{th}$ client, we draw $D_{s}(c)=\lfloor\frac{\alpha}{c+r}\rfloor+\gamma$ samples without replacement from two arbitrarily selected classes. $r$ ensures $D_{s}$ doesn’t outnumber class samples for small $c$ values. We set $\alpha=2000$, $\gamma=20$, and $r=2$.
Fig. 3 shows an example of such distributions for the MNIST dataset.
5.1.3 Defense setting
We use a CVAE conditioned on class labels. Its encoder and decoder are two-layer MLPs with a hidden layer size of 100 neurons. The number of CVAE training epochs has been fixed to 20. The KLD annealing weight was set to 0 with an increase of 0.5 after 10 epochs. The server $S$ trained its global model copy for 10 warmup epochs, then saved the activation maps from the next 10 epochs with a trigger set containing 250 samples extracted from the test set. We perform trust propagation with $\lambda=0.3$.
Figure 3: Non-IID distributions of the MNIST training samples among 20 clients, colors represent target classes
5.2 Considered Byzantine attacks
Four attacks tested and parametrized in the literature were tested [12, 24, 6]:
  • Sign-Flipping attacks: Attackers reverse the signs of their model updates by multiplying them by a factor of $-\xi$. This is akin to reversing gradient descent and leading the model in the direction of errors. Let ${\omega_{i}}$ be the weights of the model update after local training. Attackers send back ${\omega^{\prime}_{i}=-\xi\omega_{i}}$.
  • Additive Noise attacks: malicious clients add Gaussian noise to some of their model weights. Attackers send back ${\omega^{\prime}_{i}=\omega_{i}+\epsilon_{i}}$, where $\epsilon_{i}\sim\mathcal{N}(0,\sigma)$.
  • Same Value attacks: malicious clients send a constant value $c$ as weight updates: ${\omega^{\prime}_{i}=c}$.
  • Scaling attack: malicious clients perform local SGD, then simply send back a scaled version of their updates, i.e., ${\omega^{\prime}_{i}=a\omega}$ with $a>1$
Attacks’ parameters $\xi$, $\sigma$, $a$, and $c$ serve as amplification factors that increase the effect of the attack at the cost of making it more detectable For our tests to be nontrivial, we worked with non-upscaled attacks that send updates close to those of benign clients by fixing $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$. We will see that with no defense, these attacks are successful.
5.3 Experimental results
Table 1: Defense results with the custom distribution
MNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 93.69 84.92 92.78 92.19 91.91 FedCAM 93.02 92.09 93.97 11.29 92.83 FedCVAE 80.82 77.15 11.32 18.68 50.23 FedGuard 61.35 17.67 10.26 11.32 10.26 FLEDGE 93.30 88.06 92.56 92.04 14.19 NoDefense 93.61 18.34 11.32 11.32 12.00 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 50.60% 0.00% FedCVAE - 0.00% 50.00% 44.60% 1.42% FedGuard - 26.89% 22.15% 12.84% 17.46% FLEDGE - 0.00% 0.00% 0.00% 70.93% NoDefense - - - - - FedCLEAN 0.08% 0.00% 0.00% 0.00% 0.00% FedCAM 47.96% 0.00% 0.00% 47.12% 0.00% FedCVAE 46.79% 0.00% 43.43% 44.60% 1.67% FedGuard 82.01% 69.28% 75.33% 87.28% 82.51% FLEDGE 4.17% 0.00% 0.00% 0.00% 35.08% NoDefense - - - - - FshMNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 69.99 64.52 66.83 66.92 65.83 FedCAM 70.06 60.47 70.83 10.03 57.17 FedCVAE 65.02 64.93 10.07 17.92 60.30 FedGuard 17.29 20.08 10.04 10.04 10.04 FLEDGE 68.78 67.82 74.39 69.95 10.09 NoDefense 71.14 18.86 10.07 13.81 11.97 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 51.40% 0.00% FedCVAE - 0.00% 51.11% 52.05% 2.68% FedGuard - 27.43% 11.86% 24.11% 18.81% FLEDGE - 0.00% 0.00% 0.00% 70.36% NoDefense - - - - - FedCLEAN 0.68% 0.00% 0.00% 0.00% 0.00% FedCAM 47.48% 0.00% 0.00% 49.51% 0.00% FedCVAE 45.61% 0.00% 43.01% 45.20% 5.98% FedGuard 95.91% 70.90% 83.49% 73.33% 80.52% FLEDGE 3.38% 0.00% 0.00% 0.00% 27.78% NoDefense - - - - -
Table 2: Defense results with the Dirichlet distribution
MNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 95.52 95.39 95.39 95.42 95.51 FedCAM 95.71 95.59 95.57 10.29 95.44 FedCVAE 94.56 93.86 11.32 10.12 17.05 FedGuard 93.20 16.69 11.32 10.33 13.93 FLEDGE 95.58 95.61 95.82 95.31 8.59 NoDefense 95.76 15.63 11.32 10.33 9.76 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 53.15% 0.00% FedCVAE - 0.00% 56.64% 60.43% 6.54% FedGuard - 21.37% 43.48% 34.14% 45.23% FLEDGE - 0.00% 0.00% 0.00% 50.81% NoDefense - - - - - FedCLEAN 0.04% 0.00% 0.00% 0.00% 0.00% FedCAM 48.04% 0.00% 0.00% 47.97% 0.00% FedCVAE 42.82% 0.00% 49.30% 46.54% 4.90% FedGuard 55.99% 77.73% 58.64% 66.06% 53.67% FLEDGE 0.04% 0.00% 0.00% 0.00% 39.07% NoDefense - - - - - FshMNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 70.86 70.47 68.99 71.06 71.85 FedCAM 74.37 72.55 71.43 10.03 70.19 FedCVAE 68.08 68.09 10.07 17.89 66.75 FedGuard 60.03 15.92 10.07 10.07 12.13 FLEDGE 71.75 70.71 70.40 70.55 9.06 NoDefense 74.39 17.88 10.07 10.01 11.95 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 52.39% 0.00% FedCVAE - 0.00% 49.35% 56.55% 2.08% FedGuard - 34.65% 31.53% 29.27% 30.50% FLEDGE - 0.00% 0.00% 0.00% 85.71% NoDefense - - - - - FedCLEAN 0.12% 0.00% 0.00% 0.00% 0.00% FedCAM 48.72% 0.00% 0.00% 48.91% 0.00% FedCVAE 44.29% 1.08% 41.91% 45.92% 4.21% FedGuard 70.00% 66.30% 70.67% 72.55% 70.22% FLEDGE 2.59% 0.00% 0.00% 0.00% 13.44% NoDefense - - - - -
Tables 1 and 2 report the results of testing on different distributions of the MNIST and FashionMNIST datasets. We use three standard metrics to measure the effectiveness of each defense system: the global model’s accuracy on the test dataset (ACC), the attacker detection false negative (FN), which refers to the fraction of attackers that manage to get through the defense system, and benign detection false positive (FP) which represents the fraction of blocked benign clients.
For a byzantine attack, the defense requirement R1 requires FN to be 0% in order for the defense to be considered successful. The defense also needs to not disrupt the training (R2), which means ACC should be as close as possible to the benchmark no defense and no attack accuracy, and FP should be as close as possible to 0%.
As we can observe, FedCLEAN is the only defense system that achieves 0% of FN against all types of Byzantine attacks in all tested scenarios. We also see that when attackers are present, FedCLEAN does not block any benign clients (i.e FP=0%). When attackers are absent, it misclassifies less than 1% of benign clients. Additionally, across 20 tested scenarios, FedCLEAN achieves the highest test accuracy in 10 scenarios and matches benchmark accuracy in all 20. But one should note that in all 10 scenarios where FedCLEAN does not achieve the highest accuracy, it achieves 100% attacker detection and 0% or close misclassification, which suggests that the gap in accuracy is due to the random attribution of samples at the start of each experiment.
6 Conclusion & future work
We introduced FedCLEAN, an autoencoder-based defense system designed to identify malicious clients in Non-IID federated learning environments from updated models’ activation maps. With two core components: A framework for training CVAEs for an FL defense system leveraging model activation maps normalized with GeoMed, DKL annealing, and a realistic trigger set requirement. And a dynamic one-parameter client selection algorithm that minimizes the exclusion of benign clients.
We demonstrate that, in image classification benchmarks, FedCLEAN is robust against Byzantine attacks under a Non-IID client data distribution and achieves top performance in benign client missclassification tests. Future work will focus on broader machine learning applications, backdoor attacks, as well as improving our method by enhancing CVAE training through adaptive regularization, and implementing a fully non-parametric trust propagation algorithm.
6.0.1 Acknowledgements
This work was partly supported by the Inserm industrial chair CYBAILE and the PEPR digital health SSF-ML-DH project, ANR-13-LAB2-0007-01

### v2

@@
-FedCLEAN: byzantine defense by CLustering Errors of Activation maps in Non-IID federated learning environments
+FL-CLEANER: byzantine and backdoor defense by CLustering Errors of Activation maps in Non-iid fedErated leaRning
 Abstract
-Federated Learning (FL) enables clients to collaboratively train a global model using their local datasets while reinforcing data privacy.
-However, FL is susceptible to poisoning attacks.
+Federated Learning (FL) enables clients to collaboratively train a global model using their local datasets while reinforcing data privacy, but it is prone to poisoning attacks.
 Existing defense mechanisms assume that clients’ data are independent and identically distributed (IID), making them ineffective in real-world applications where data are non-IID.
-This paper presents FedCLEAN, the first defense capable of filtering attackers’ model updates in a non-IID FL environment.
-The originality of FedCLEAN is twofold.
+This paper presents FL-CLEANER, the first defense capable of filtering both byzantine and backdoor attackers’ model updates in a non-IID FL environment.
+The originality of FL-CLEANER is twofold.
 First, it relies on a client confidence score derived from the reconstruction errors of each client’s model activation maps for a given trigger set, with reconstruction errors obtained by means of a Conditional Variational Autoencoder trained according to a novel server-side strategy.
-Second, we propose an ad-hoc trust propagation algorithm based on client scores, which allows building a cluster of benign clients while flagging potential attackers.
-Experimental results on the datasets MNIST and FashionMNIST demonstrate the robustness of FedCLEAN against Byzantine attackers in non-IID scenarios and a close-to-zero benign client misclassification rate, even in the absence of an attack.
-Keywords:
-Poisoning attacks Byzantine attacks Non-IID Federated Learning Conditional variational autoencoders Clustering
+Second, it uses an original ad-hoc trust propagation algorithm we propose.
+Based on previous client scores, it allows building a cluster of benign clients while flagging potential attackers.
+Experimental results on the datasets MNIST and FashionMNIST demonstrate the efficiency of FL-CLEANER against Byzantine attackers as well as to some state-of-the-art backdoors in non-IID scenarios; it achieves a close-to-zero (<1%) benign client misclassification rate, even in the absence of an attack, and achieves strong performance compared to state of the art defenses.
+Keywords Poisoning attacks, Byzantine attacks, Backdoor attacks, Non-IID Federated Learning, Conditional variational autoencoders, Clustering.
 1 Introduction
@@
 At first, a central server initiates the global model and sends it to the clients so that they can train it locally on their private data.
-Model updates are sent back to the server, which aggregates them into a new global model.
+Model updates, in the form of weights, are sent back to the server which aggregates them into a new global model.
 These two steps are repeated until the global model converges.
@@
 However, even though FL systems preserve privacy, their decentralized nature makes them vulnerable to poisoning attacks, where one or several malicious clients try to corrupt or prevent the training session by sending malicious updates [7].
-Some defenses try to mitigate these attacks using secure aggregation algorithms with the idea of reducing the impact of bad updates [41, 8, 26].
-Nevertheless, these solutions remain limited, being, for instance, not resistant to attacks that can destroy the global model after just one round of a successful attack, like Byzantine attacks [12].
+Some defenses try to mitigate the impact of these attacks using secure aggregation algorithms with the idea of reducing the impact of bad updates [8, 9, 10].
+Nevertheless, these solutions remain limited.
+They are, for instance, not resistant to attacks able to destroy the global model after just one successful FL round of attack, known as Byzantine attacks [11].
 This has led to the proposal of FL defense systems devoted to the detection of malicious updates before the aggregation stage.
@@
 However, such systems do not perform well in non-IID contexts [17], which are more likely encountered in real-life scenarios.
-In the following, we introduce these solutions and their limitations, then present our contribution: FedCLEAN, the first FL defense system adapted to non-IID contexts.
+In the following, we introduce these solutions pointing out their limitations before then presenting our contribution: FL-CLEANER, the first FL defense system adapted to non-IID contexts.
+Before concluding, we will also experimentally show that FL-CLEANER: i) effectively blocks all byzantine attacks in a non-IID setting, all while having near zero false positives; ii) mitigates the effects of backdoor attacks.
+Figure 1: A common online FL scenario with a central server
 1.1 Related Works
-As stated above, the first defenses against malicious updates were the development of robust aggregation algorithms.
+As stated above, the first defenses against malicious clients’ updates were the development of robust aggregation algorithms.
 FedAvg [8] is one of the earliest and most commonly used aggregation methods in FL.
@@
 Clients’ dataset sizes are used as weights.
-Doing so dilutes the effect of sparse attackers but is not sufficient when their number increases.
-It is also not robust to FL attacks that amplify updates to dominate the global average [48], and ineffective in non-IID scenarios where an attacker has a larger dataset compared to benign clients.
+Doing so dilutes the effect of attackers when they are few in number but is not sufficient when their number increases.
+It is also not robust to FL attacks that amplify updates to dominate the global average [18], and ineffective in non-IID scenarios especially when an attacker has a larger dataset compared to benign clients.
 GeoMed [10] uses the geometric median of clients’ updates as a global update.
@@
 [23] were the first to propose such a defense.
-It trains a copy of the global model on the test dataset on the server side, while training at the same time a variational autoencoder (VAE) on surrogate vectors randomly sampled from the weights of this model at each epoch.
+It trains a copy of the global model on the test dataset on the server side, while training at the same time a variational autoencoder (VAE) on surrogate vectors (randomly sampled weights from the model) at each epoch.
 Although this solution performs better than aggregation-based defenses under Byzantine attacks, it requires training a VAE on an entire server-side dataset to achieve good results.
@@
 FedCVAE shows some defense results in a non-IID scenario but only against attacks exaggerated in magnitude and thus fairly easy to detect.
+They also provide results on simple backdoor attacks.
 More recently, FedCAM [16], a CVAE anomaly detection-based approach, achieves the latest state-of-the-art results against Byzantine attacks in IID scenarios.
@@
 They have shown promising results on anomaly detection without needing autoencoders.
-Among recent ones, we can cite DeepSight [46].
-It uses a model inspection technique and a clustering scheme to create clusters of updates, and to pick out those with the best scores.
-DeepSight successfully stops backdoor attackers [27], even in the non-IID case, by selecting groups with minimal attacker influence.
-However, it does not work for Byzantine attacks where a single successful breach is sufficient to make the global model convergence fail.
-More recently, FLEDGE [10] used a combination of cosine similarity and clustering to filter out attacks.
-It provides results against Byzantine attacks but only a fraction of their data is non-IID.
+Among recent ones, we can cite FLEDGE [14] which used a combination of cosine similarity and Kernel Density Estimation to filter out attacks.
+It focuses on backdoor attacks but provides results against Byzantine attacks but only a fraction of their data is non-IID.
+Finally, Sharma et al.
+[25] recently proposed a defense based on graph representation combined with Graph Attention Networks.
+Beyond the fact they require pre-training a model on both attacker and benign client examples, its authors provide only results for one byzantine attack, being sign-flipping attack.
+In addition to these defenses, several others exist that specifically target backdoor attacks [26].
+Defenses such as FoolsGold [27], DeepSight [15], FLAME [28] and FLDetector [29]rely on filtering out attackers by finding clients that present similar behavior.
+Other propositions try to remove the effect of the backdoor after the aggregation step, such as [30] and [31].
+However, these propositions have not been tested or proven effective against byzantine attacks.
 1.2 Contribution
-In this work, we propose FedCLEAN, an extension of FedCAM, a solution introduced by Bellafqira et al.
-[6].
-Similar to FedCAM, FedCLEAN analyzes the GeoMed normalized CVAE reconstruction errors of clients’ activation maps for given inputs of a trigger set.
+In this work, we propose FL-CLEANER, an extension of FedCAM, introduced in [16], which was only devoted to byzantine attack detection in IID settings.
+Similarly to FedCAM, FL-CLEANER analyzes the GeoMed normalized CVAE reconstruction errors of clients’ activation maps for given inputs of a trigger set.
 Clients with high reconstruction errors are rejected based on an anomaly detection strategy.
-While FedCAM achieves top byzantine robustness compared to concurrent methods in IID scenarios [6], it fails to achieve good results in the context of non-IID data, as we will see in Section 5.
-We identified two main issues causing this failure, which we address with FedCLEAN.
+While FedCAM achieves very good byzantine robustness compared to concurrent methods in IID scenarios, it fails to achieve good results in the context of non-IID data, as we will see in Section 5.
+Two main issues cause this failure.
 First, the CVAE training method proposed in FedCAM has poor generalization in non-IID contexts, as client updates are more heterogeneous.
 Specifically, clients with the most unique data often fall out of distribution for CVAE and end up with similar error scores to attackers.
-To overcome this issue, FedCLEAN employs a different CVAE training process that accounts for the temporal evolution of FL training with warmup epochs and better captures the statistical properties of benign updates using an annealing technique and layer selection.
+To overcome this issue, FL-CLEANER employs a different CVAE training process that accounts for the temporal evolution of FL training with warmup epochs and better captures the statistical properties of benign updates using an annealing technique and layer selection.
 This ensures that all benign client activation maps, even outliers, are correctly reconstructed.
 Second, FedCAM relies on the mean of reconstruction errors as a dynamic error threshold.
-This approach risks allowing the selection of some attackers’ updates for the aggregation step if the average is high enough.
+This approach can allow the selection of some attackers’ updates for the aggregation step if the average is high enough.
 It is important to note that retaining even one Byzantine attacker in a round is sufficient to destroy the global model.
-Moreover, this strategy also impacts the convergence or accuracy of the global model in the absence of attackers, as it eliminates around half of the participants in each round.
-In the non-IID case, clients are less redundant compared to IID data.
-To address this problem, FedCLEAN introduces a clustering algorithm that iteratively aggregates and concentrates benign candidates into a trusted cluster, enabling a more refined selection process.
+Moreover, this strategy eliminates around half of the participants in each round.
+In the non-IID case where clients are less redundant compared to IID scenarios, this impacts the convergence and accuracy of the global model.
+To address this problem, FL-CLEANER introduces a clustering algorithm that iteratively aggregates and concentrates benign candidates into a trusted cluster, enabling a more refined selection process.
 The rest of this paper is organized as follows.
@@
 Section 3 presents the federated context we considered and its threat model.
-Section 4 introduces our proposal, FedCLEAN, followed by Section 5, which details our benchmark and provides an analysis of the experimental results.
+Section 4 introduces our proposal, FL-CLEANER, followed by Section 5, which details our benchmark and provides an analysis of the experimental results.
 Finally, we conclude the paper in Section 6.
@@
 In this work, we considered an FL setting where $N$ clients $\{C_{k}\}_{k=1..N}$ with their datasets $\{D_{k}\}_{k=1..N}$ participate in a training session orchestrated by a central server $S$ to train a global model $M_{G}$.
-During each FL round $t$, $S$ scatters the current global model $M_{G}^{t}$ to a subset of $U$ randomly selected clients.
-Each client updates the received model on its local dataset for a number of local epochs and sends back the updated model to $S$.
+During FL round $t$, $S$ scatters the current global model $M_{G}^{t}$ to a subset of $U$ randomly selected clients.
+Each client updates the received model on its local dataset and sends back the updated model to $S$.
 The server then uses an aggregation algorithm such as FedAvg [8] to get an updated version of the global model $M_{G}^{t+1}$ from the clients’ updated models $\{M_{u}^{t}\}_{u=1..U}$.
@@
   • Quantity skew: Where the amount of local data may be unbalanced across clients.
-In practice, feature and quality skew are intrinsic properties of data and are hard to simulate for an experiment if the dataset does not provide a natural partition.
-So, most works that address Non-IID FL use a combination of label and quantity skew by choosing a distribution of data over clients that ensures heterogeneity.
-2.1.3 Attacks in Federated Learning
+Non-IID data hypothesis is considered a challenge for FL defenses for two main reasons.
+First, a well known problem in FL literature [36, 32] is that it is harder to make the global model converge in this setting, since clients with different data often send contradicting updates.
+From a defense point of view, this means that it is a risk to misclassify honest clients, requiring precise clients’ udpates filtering.
+Second, defenses that rely on detecting attackers often look for statistical anomalies to spot these attackers.
+This is more easy in a IID data context where clients share similar data distribution and thus present similar statistical properties.
+With non-IID data, clients naturally have different characteristic, and it becomes more difficult to differentiate attackers from natural outliers.
+In practice, feature and quality skew are intrinsic properties of data and are hard to simulate experimentally if the dataset does not provide a natural partition.
+So, most works that address Non-IID FL use a combination of label and quantity skews by choosing a distribution of data over clients that ensures heterogeneity.
+2.1.3 Attacks on Federated Learning
 Federated Learning is known to be vulnerable to several attacks [32].
-During training, these attacks involve an adversary that compromises one or several clients and tries to use them to disrupt the global model.
-Herein, two main approaches can be distinguished: "model poisoning", where the attacker can directly manipulate the compromised clients’ models and send malicious updates to the server; and "data poisoning", where the attacker indirectly manipulates learning by corrupting the clients’ local datasets.
+These attacks operate at training time and involve an adversary that compromises one or several clients and tries to use them to disrupt the global model.
+Two main approaches can be distinguished: "model poisoning", where the attacker can directly manipulate the compromised clients’ models and send malicious updates to the server; and "data poisoning", where the attacker indirectly manipulates learning by corrupting the clients’ local datasets.
 Based on the attacker’s objective, these attacks can further be divided into two sub-classes:
   • Untargeted poisoning attacks - where the attacker’s goal is to drop the global model accuracy on any future inputs by preventing its convergence.
-It is generally very hard for a machine learning model to recover from a bad convergence, and one attack in an unprotected FL round is often sufficient for an untargeted poisoning attack to succeed.
+It is generally very hard for a machine learning model to recover from a bad convergence, and one attack in an unprotected FL round is often sufficient for an untargeted poisoning attack to succeed [9].
 Such attacks are also referred to as "Byzantine attacks".
@@
 The most common form of backdoor attacks involves inserting a trigger in the training dataset and training the model to output a specific response when the trigger data are given as an input.
-It is important to notice that making a successful backdoor attack is very hard in Non-IID scenarios [27].
-Controlling just some of the FL clients with non-IID data, the attacker generally has access only to a subpart of the training data, which in some cases is not sufficiently representative for injecting a robust backdoor to benign clients’ updates [39].
-Moreover, the backdoor updates are easily overwritten by benign clients’ ones, as shown in [39].
-This is the reason why in this work, we decided to focus on defending against Byzantine attacks, which currently pose a greater threat in FL.
+[37].
+Through the control of just some of the FL clients, the attacker may have insufficiently representative data because of the Non-IID context [38], making the backdoor updates easily overwritten by future training rounds, as shown in [38]
 2.2 Conditional Variational Autoencoders (CVAE)
@@
 The objective of an autoencoder is to minimize the reconstruction error (RE) between the initial and reconstructed vectors.
-They have been historically used for dimensionality reduction [4, 30], model pre-training [20, 7], and generative tasks [52].
+They have been historically been used for dimensionality reduction [40, 39], model pre-training [41, 42], and generative tasks [43].
 An important property of AEs is that, when applied to out-of-distribution vectors that deviate from the original training distribution, these deviations are amplified in the latent space, and much higher reconstruction errors are observed.
@@
 Conditional-VAEs (CVAEs) [47] are an extension of VAEs that leverages conditional probabilities to guide the encoding and decoding process following a provided condition.
-In practice, what differentiates a CVAE from a traditional VAE is the addition of a condition vector $y$ as an input of both the encoder and decoder (see Fig.
-1).
-Thus, CVAEs enhance the capabilities of VAEs by making them compatible with supervised learning and allowing them to work in a task-guided manner [16].
-They have been extensively used for a number of applications, including anomaly detection and detecting attackers in federated learning [24, 6, 11].
+In practice, a CVAE is a VAE with a condition vector $y$ as additional input for both the encoder and decoder (see Fig.
+2).
+CVAEs have been extensively used for a number of applications, including anomaly detection and detecting attackers in federated learning [12, 16, 13].
 2.3 GeoMed and Activation Maps
@@
 This has led to its extensive use in Byzantine-robust federated learning as a robust aggregation method [48] or to detect attackers [12].
-In a typical deep learning model, the data is processed through feed-forward layers.
-Each hidden layer takes in the previous layer’s output, applies a linear transformation parametrized by a set of weights (e.g., a convolution or weighted sum), and then applies a nonlinearity function.
+In a typical deep learning model, data are processed through feed-forward layers.
+Each hidden layer takes in the previous layer’s output, applies a linear transformation parameterized by a set of weights (e.g., a convolution or weighted sum), and then applies a nonlinearity function.
 The output of the layer’s weighted transformation is referred to as the activations of that layer.
-"Activation maps", or the set of activation maps extracted from a model for a given input, are characteristic of a model and the way it processes data.
+"Activation maps", or the set of activations extracted from a model for a given input, are characteristic of a model and the way it processes data.
 They have been used for applications such as explainability [49] and model watermarking [50, 51, 52].
@@
 3.1.1 Adversary’s Objective
-In this work, we consider a "Byzantine" adversary $A$, who wishes to carry out Byzantine attack during the training phase of an FL model.
+In this work, we consider a "Byzantine/backdoor" adversary $A$, who wishes to carry out Byzantine/backdoor attack during the training phase of an FL model.
 To do so, $A$ controls a number of the FL clients who submit poisoned model updates to $S$ during the FL process with the objective of disturbing the convergence of the global model.
-Formally, $A$ attempts to fulfill the following objectives:
-  • O1: If the attack succeeds, the global model’s accuracy on test data should collapse.
+Formally, for Byzantine attacks, $A$ attempts to fulfill the following objectives:
+  • O1: If the attack succeeds, the global model should have bad accuracy on test data.
   • O2: The attacked model should not be able to recover from the accuracy drop by pursuing further training in the absence of the attacker.
+And, for a backdoor to be successful :
+  • O3: Achieve a good attacker success rate (ASR), which corresponds to the ratio of test samples with the backdoor trigger that the model successfully classifies in the target class.
 3.1.2 Adversary’s Capabilities
-As in [3, 46, 22, 11, 24, 10], we consider a strong adversary who has full control over $m<N/2$malicious" clients.
-$A$ has full access to these clients’ datasets, models, and communication channels.
+As in [54, 15, 55, 13, 12, 14], we consider a strong adversary who has full control over $m<N/2$ "malicious" clients; $N$ being the total number of clients.
+$A$ has full access to datasets, models, and communication channels of these clients.
 In particular, $A$ can conduct training using local datasets when clients are selected by the server, freely modify the weights of the resulting model updates, and send back modified updates to the server $S$ for aggregation.
@@
 3.2 Security Requirements
-To effectively defend against a Byzantine attacker, it is required that a filtering-based defense system $D$ satisfies the following requirements [46]:
-  • R1: Block all malicious clients.
-As discussed above, the Byzantine adversary $A$ is able to break the model given one single successful round of attack.
-As a result, in order to be robust to such an adversary, $D$ must detect all malicious clients updates at any round.
+To effectively defend against a Byzantine or backdoor attacker, it is required that a filtering-based defense system $D$ satisfies the following requirements [15]:
+  • R1: Block all malicious byzantine clients.
+As discussed above, the Byzantine adversary is able to break the model given one single successful round of attack.
+As a result, in order to be robust to such an adversary, $D$ must detect all of these malicious client updates at any round.
   • R2: Do not disrupt training.
@@
 In other words, $D$ must have high precision and not misclassify benign clients.
-It must also preserve the test accuracy achieved in the no-attack and no-defense scenario.
-We will see that our proposal responds to all these objectives and security requirements.
-4 Proposed defense: FedCLEAN
-4.1 FedCLEAN Principle
-The idea behind FedCLEAN is to analyze the normalized activation maps (NAMs), obtained by subtracting the GeoMed value from activation maps of one or several layers of $M_{G}$ for some given inputs referred as a trigger set $T_{s}=(X_{T},Y_{T})$, where $X_{T}$ and $Y_{T}$ are the set of samples and their labels, respectively.
-We assume that the NAMs obtained with malicious model updates will correspond to outliers that can be identified by means of an CVAE and GeoMed based anomaly detection strategy.
-More clearly, before the FL process begins : FedCLEAN : i) makes a copy of the initial global model; ii) trains the copy locally for several epochs on a small dataset called the "trigger set"; iii) extracts the NAMs associated with the trigger set from different epochs of training; iv) trains a CVAE to reconstruct these NAMs, using their corresponding labels $Y_{T}$ as a condition.
-The training process is detailed in 4.2.
+It must also preserve the test accuracy achieved in the baseline no-attack and no-defense scenario.
+  • R3: Lower backdoor attacker success rate. In the presence of the defense, the backdoor accuracy measured on the test set should be significantly lower.
+We will see in the sequel that our proposal fulfills these three security requirements.
+4 Proposed defense: FL-CLEANER
+4.1 FL-CLEANER Basic Principles
+The idea behind FL-CLEANER is to analyze the normalized activation maps (NAMs), obtained by subtracting the GeoMed value from activation maps of one or several layers of $M_{G}$ for some given inputs referred as a trigger set $T_{s}=(X_{T},Y_{T})$, where $X_{T}$ and $Y_{T}$ are the set of samples and their labels, respectively.
+We assume that NAMs obtained with malicious model updates will correspond to outliers that can be identified by means of a CVAE and GeoMed based anomaly detection strategy.
+In a simplified view, before the FL process begins, FL-CLEANER :
+  1. 1.
+    Makes a copy of the initial global model.
+  2. 2.
+    Trains this copy locally for several epochs on a "trigger set" that is a very small dataset.
+  3. 3.
+    Extracts NAMs associated with the trigger set from different epochs of training
+  4. 4.
+    Trains a CVAE to reconstruct these NAMs, using their corresponding labels $Y_{T}$ as a condition. The training process is detailed in 4.2.
 By doing so, the CVAE learns the latent distribution of NAMs associated with each class.
-Once trained, the CVAE is used by $S$ at each epoch to reconstruct for each client updates the NAMs corresponding to the trigger set.
-Malicious NAMs, which fall out of the learned distribution, result in high reconstruction errors.
-To discriminate malicious NAMs from benign ones, FedCLEAN makes use of a clustering procedure relying on an ad-hoc trust propagation algorithm.
-We discuss it in subsection 4.3.
-To provide a formal description of FedCLEAN, let us consider a classification task and $N$ clients $\{C_{k}\}_{k=1..N}$. FedCLEAN operates according to the following steps :
+Once trained, $S$ uses CVAE at each epoch to reconstruct NAMs corresponding to the trigger set for each client updated model $S$ received.
+Malicious client NAMs, which fall out of the learned distribution, result in high reconstruction errors.
+To discriminate malicious NAMs from benign ones, FL-CLEANER makes use of a clustering procedure relying on an ad-hoc trust propagation algorithm.
+We present and discuss it in subsection 4.3.
+To go in further details and provide a formal description of FL-CLEANER, let us consider a FL scenario with a classification task and $N$ clients $\{C_{k}\}_{k=1..N}$.
+FL-CLEANER operates according to the following steps:
   1. 1.
-    FL and FedCLEAN Initialization
+    FL and FL-CLEANER Initialization ($t=0$)
     1. (a)
-      $S$ initializes the global FL model $M_{G}^{0}$ according to a common process. By assigning, for instance, random values to its weights.
+      $S$ initializes the global FL model $M_{G}^{0}$ in a classic way assigning, for instance, random values to its weights.
     2. (b)
@@
     3. (c)
-      $S$ then trains the CVAE.
+      $S$ then trains a CVAE model.
 To do so, it first extracts the activation maps $A_{S}$ of one or more layers from $M_{G}^{{}^{\prime}0}$ feeding it with the trigger set $T_{s}$.
-Then, it normalizes them using GeoMed such as: $\hat{A}_{S}=\sigma(A_{S}-\text{GeoMed})$ (where $\sigma$ is the sigmoid function).
-$S$ finally trains the CVAE on $\hat{A}_{S}$.
+Then, it normalizes them using GeoMed such as: $\hat{A}_{S}=\sigma(A_{S}-\text{GeoMed})$, with $\sigma$ is the sigmoid function.
+$S$ finally trains CVAE on $\hat{A}_{S}$.
     4. (d)
-      $S$ then sends $M_{G}^{0}$ to $U$ randomly selected clients ($U<N$), to initiate the first FL training round.
+      $S$ initiates the first FL training round by sending $M_{G}^{0}$ to $U$ randomly selected clients ($U<N$),.
   2. 2.
-    FL training process for each round $t>1$
+    Round $t\geq 1$ of the FL training process
     1. (a)
@@
     2. (b)
-      Secure FedCLEAN selection
+      Secure FL-CLEANER selection
       1. i.
-        Inspection the activation maps for each client $u$: $S$ extracts the activation maps $A_{u}^{t}$ obtained by feeding the trigger set $T_{s}$ into $M^{t}_{u}$.
+        Activation map inspection for each client $u$ : $S$ extracts the activation maps $A_{u}^{t}$ obtained by feeding the trigger set $T_{s}$ into $M^{t}_{u}$.
       2. ii.
         Scoring:
-        • $S$ computes the activation maps’ GeoMed: $\text{GeoMed}(A_{1}^{t},...,A_{U}^{t})$ and normalizes the activation maps of each client: $\hat{A}_{u}^{t}=\sigma(A_{u}^{t}-\text{GeoMed})$.
-        • $S$ feeds its CVAE with $\hat{A}_{u}^{t}$ conditioned on $Y_{t}$ and assigns to each client a reconstruction error score $\epsilon_{u}^{t}=\overline{MSE}(\hat{A}_{u}^{t},\tilde{A}_{u}^{t})$ where $\tilde{A}_{u}^{t}$ is the CVAE reconstruction of $\hat{A}_{u}^{t}$, and $MSE$ the mean square error function.
+        • $S$ computes $\text{GeoMed}(A_{1}^{t},...,A_{U}^{t})$, that is the GeoMed centroid of all activation maps.
+Then it normalizes the activation maps of each client: $\hat{A}_{u}^{t}=\sigma(A_{u}^{t}-\text{GeoMed})$.
+        • $S$ feeds its CVAE with $\hat{A}_{u}^{t}$ conditioned on $Y_{t}$ , the label, and assigns to each client a reconstruction error score $\epsilon_{u}^{t}=\overline{MSE}(\hat{A}_{u}^{t},\tilde{A}_{u}^{t})$ where $\tilde{A}_{u}^{t}$ is the CVAE reconstruction of $\hat{A}_{u}^{t}$, and $MSE$ the mean square error function.
     3. (c)
-      Filtering: Given the pairs $(M_{u}^{t},\epsilon_{u}^{t})$, $S$ performs our trust propagation clustering to filter out clients that are considered malicious and discards their updates.
+      Filtering: Given the pairs $(M_{u}^{t},\epsilon_{u}^{t})$, $S$ performs our trust propagation clustering (see next Section) to filter out clients that are considered malicious and discards their updates.
     4. (d)
       FL Aggregation: Finally, $S$ aggregates updates from the remaining clients using FedAvg to generate $M_{G}^{t}$.
-As exposed, the performance of FedCLEAN relies on two main components a CVAE and a trust propagation clustering algorithm used to discriminate malicious from benign updates.
+As exposed, the performance of FL-CLEANER to discriminate malicious from benign updates rely on two main components: a CVAE model and a trust propagation clustering algorithm.
 We come back to how they work in the following sections.
-4.2 CVAE training constraints
-As stated above, the CVAE model is trained by the server $S$ considering a simulated a benign model $M_{G}^{{}^{\prime}0}$ based on the server trigger set $T_{s}$.
-One of the constraints we assumed in building FedCLEAN is that $T_{s}$ is of very small size compared to the client data.
-To train the CVAE well, one thus needs to augment the training set.
-The solution we adopted is to use the activation maps (AMs) computed at different rounds of the training of $M_{G}^{{}^{\prime}0}$.
-But by doing so, we have to take into account the temporal drift of $M_{G}^{{}^{\prime}0}$ activation maps that impact CVAE generalization capabilities.
-In fact, AMs are more heavily affected by noise and very different at the very beginning of the training of $M_{G}^{{}^{\prime}0}$ than after a certain number of epochs.
-To overcome this issue, and as illustrated in Fig.
-2 the CVAE training set is constituted of activation maps collected after some warmup epochs of the training of $M_{G}^{{}^{\prime}0}$.
-At each epoch, AMs are GeoMed normalized so as to better control AMs amplitude differences along the training of $M_{G}^{{}^{\prime}0}$.
-Indeed, with a GeoMed recomputed at each round, we have found that the distribution of the deviations from GeoMed remains stable.
-Anyway, even with a small trigger set, this technique allows to constitute a training set that captures the distribution of benign NAMs assuming certain hypotheses.
+4.2 CVAE training
+As stated above, the CVAE model is trained by the server $S$ considering a simulated benign model $M_{G}^{{}^{\prime}0}$ based on the server trigger set $T_{s}$.
+In this work we assume that $T_{s}$ is of very small size compared to the client data.
+This ensures the real-life applicability of our solution as servers may not have access to the large datasets other methods in literature require.
+Moreover, servers may have some data of their own or access to trusted clients or can form such datasets from open-source data.
+The influence of the trigger set size and content is not included in our study due to space limitations.
+However, our experiments showed that it is possible to achieve good defense results against Byzantine attackers with as little as 250 images randomly sampled from the server test set and 500 images for backdoor attacks.
+To augment the number of training samples available to the CVAE with a small dataset, the solution we adopted is to use the activation maps (AMs) computed at different rounds of the training of $M_{G}^{{}^{\prime}0}$.
+AMs are more heavily affected by noise and vary a lot from one epoch to another at the very beginning of the training of $M_{G}^{{}^{\prime}0}$.
+To account for this, as whown in in Fig.
+3, the CVAE training set is constituted of activation maps collected after some warmup epochs of the training of $M_{G}^{{}^{\prime}0}$.
+Additionnally, AMs are GeoMed normalized into NAMS at each epoch so as to better control AMs amplitude differences along the training of $M_{G}^{{}^{\prime}0}$.
+Doing so also helps to take into account the temporal drift of $M_{G}^{{}^{\prime}0}$ activation maps.
+Indeed, the distribution of the deviations from GeoMed recomputed at each iteration remains stable throught the epochs.
+That is not the case of the distribution of the AMs which drifts over time.
+We will see that even with a small trigger set, these technique allows to constitute a training set that well captures the distribution of benign NAMs, and ensures good CVAE generalization capabilities.
 Figure 3: Construction of the CVAE training set.
-A second issue we had to face when training the CVAE is that it is sensitive to "posterior collapse" [29, 9, 14], which causes it to converge to minima that ignore training inputs.
-More clearly, the traditional training of a CVAE, consists in maximizing the Evidence Lower Bound (ELBO) in the following loss:
+A second issue to consider is that CVAE training is sensitive to "posterior collapse" [56, 57, 58], causing it to converge to minima that ignore training inputs.
+Indeed, the traditional training of a CVAE, consists in maximizing the Evidence Lower Bound (ELBO) in the following loss:
 $\mathcal{L}(x,y;\theta,\phi)=-\text{ELBO}(x,y;\theta,\phi)\\ =-(\mathbb{E}[\log p_{\theta}(x|z,y)]-\frac{1}{2}[\sigma_{\phi}(x)^{2}+\mu_{\phi}(x)^{2}-1-log(\sigma_{\phi}(x)^{2}))\\ =\text{MSE}(g_{\theta}(z,y),x)+D_{\text{KL}}(q_{\phi}(z|x)\|p(z))$ (1)
-where: $g_{\theta}$ is the CVAE decoder; $q_{\phi}$ is the latent space distribution given by the CVAE encoder; $z$ is a latent representation sampled from this distribution; and, $(x,y)$ the CVAE input sample and its label; $D_{KL}$ is the Kullback-Leibler divergence measuring the divergence with a prior distribution $p$, in our case a standard normal distribution.
-As seen, this loss is constituted of a first term to minimize the CVAE reconstruction error and of a second to enforce a Gaussian structure for the latent space representation.
-In our context, NAMs give rise to numerically very small MSE values, so the optimizer focuses on minimizing the second term, leading to the so called "posterior collapse" issue.
-To solve this issue, it has been proposed to add a regularization weight to the divergence term in the loss function [29].
-In our work, we adopted a KLD annealing approach with an adaptive regularization weight $\beta^{t}$ like [9]
+where: $g_{\theta}$ is the CVAE decoder; $q_{\phi}$ is the latent space distribution given by the CVAE encoder; $z$ is a latent representation sampled from this distribution; and, $(x,y)$ the CVAE input sample and its label; $D_{KL}$ is the Kullback-Leibler divergence measuring the divergence with a standard normal prior distribution $p$.
+As exposed, the first loss term minimizes the CVAE reconstruction error and the second enforces a Gaussian structure for the latent space representation.
+In our context, NAMs give rise to numerically very small reconstruction error values making the optimizer to focus on minimizing the second term: the so called "posterior collapse" issue.
+To solve this issue, we add a regularization weight to the divergence term in the loss function [56].
+In our work, we adopted a KLD annealing approach with an adaptive regularization weight $\beta^{t}$ as in [57]
 $\mathcal{L}_{CVAE}=MSE+\beta^{t}D_{KL}$ (2)
-where, $t$ is the current training epoch.
-CVAE starts to train with $\beta^{0}=0$ so that it learns to produce meaningful representations of the feature space and make predictions based on such representations.
-Then,$\beta$ is gradually increased by a fixed increase to enforce the structure of the latent space.
+where, $t$ is the current training epoch. CVAE starts its training with $\beta^{0}=0$ Then, $\beta$ is gradually increased by a fixed increase to enforce the structure of the latent space.
 4.3 Clustering using trust propagation
 A commonplace practice in CVAE anomaly detection and FL defenses [16, 12, 23] is to threshold error values to decide of an anomaly.
-Usually, this is the average reconstruction error of all clients.
-However because of this, in the absence of an attacker, nearly half of the clients even if they are all benigns are wrongly flagged as malicious.
+Usually, this is the average reconstruction error of all clients [16, 12].
+A result of this is that in the absence of an attacker almost half of the clients are wrongly flagged as malicious, even if they are benign.
 Well-adjusted static thresholds could allow for better filtering.
@@
 However, in a Non-IID scenario, real client reconstruction error values cannot be deduced from the sample training scenario.
-We thus opted for a simple clustering algorithm instead.
-To avoid systematically excluding benign clients, our clustering algorithm is able to automatically output either one or two clusters depending on whether malicious updates are present.
-This is a challenge as deciding the number of clusters based on data properties is an active problem in cluster analysis [21], and most algorithms require fixing the number of clusters beforehand as a user-defined parameter [21].
-In addition, clustering algorithms vary widely in terms of "parameter reliance", i.e.
-the dependence on user-defined parameters and the tolerance to different ranges of such parameters.
+We opted for a simple clustering algorithm instead.
+Our clustering algorithm automatically outputs either one or two clusters depending on whether malicious updates are present.
+Notice that deciding the number of clusters based on data properties is an active problem in cluster analysis [59].
+Indeed, most algorithms require fixing the number of clusters beforehand as a user-defined parameter [59] with clustering performance that varies widely in terms of "parameter reliance", i.e.
+parameter tolerance ranges.
 Our proposal has a single parameter that is interpretable (i.e.
-not rely on guessing).
-It sequentially retrieves all benign clients participating in training, through what we call a "trust propagation" algorithm.
-It relies on a single relative parameter, which is a fraction in $[0,1]$ that can be selected in a data-agnostic manner, which it then uses to dynamically compute the distance measures it needs to perform clustering.
-Our idea relies on the fact that when reconstructing benign client activation maps, the measured reconstruction error is a residual from the randomness and uncertainty within CVAE.
-More clearly, when reconstructing out-of-distribution attacker NAMs, the reconstructed vector is an arbitrary response.
-Thus, the reconstruction error we measure is akin to measuring error to a random vector.
-As a result, assuming we are working with an autoencoder that has properly fitted benign NAMs distributions, we expect the error from the reconstruction of benign NAMs to be of a lower order of magnitude than those from anomalous NAMs.
-In other words: in the reconstruction error space, there is an identifiable relative error gap between the two.
-Our algorithm works as follows.
-It sorts the client reconstruction errors, and selects the lowest reconstruction error client as the core of the benign client group.
-Then, in the reconstruction error space, if a client is within a certain distance $\delta$ of the last accepted client, he is added to the group.
-The iterations stop when the next client is at $d>\delta$ distance from the current client.
-We assume in that case that the gap between error values corresponds to that between attackers and benigns, and block all remaining clients.
-$\delta$ is computed dynamically as a proportion of the size of the error space for the current round, as $\delta=\lambda(max(\epsilon_{u}^{t})-min(\epsilon_{u}^{t}))_{0<u<U}$, where $\lambda\in[0,1]$ is a hyper-parameter fixed at the start of the defense system.
-The full pseudo-code for the trust propagation algorithm is detailed in Alg.
-1.
+not rely on guessing) and fixes the number of clusters automatically.
+The basic idea of our solution relies on the fact that for benign client’s activation maps the measured reconstruction error is a residual from the randomness and uncertainty within CVAE and that the reconstructed vector in the case of an out-of-distribution attacker NAMs is an arbitrary response akin to measuring error to a random vector.
+Assuming that the CVAE has properly fitted benign NAMs distribution, expected reconstruction errors of benign NAMs are of a lower order of magnitude than for anomalous NAMs.
+In other words: in the reconstruction error space, there is an identifiable relative error gap between the two, as exemplified in Figure 4 in the case of a sign flipping attack (sign inversion of mdoel weights).
+Based on this, we propose a "trust propagation" algorithm the principle of which consists in sequentially retrieving all benign clients participating in training.
+First, it sorts the client reconstruction errors, and selects as the core of the benign client cluster the one with the lowest error score.
+Then, in the reconstruction error space, if a client is within a certain distance $\delta$ of the last accepted client, it is added to the group.
+The iterations stop when the distance $d$ of the next client from the current client is such as $d>\delta$.
+This gap between error values corresponds to the frontier between attackers and benigns.
+All the remaining clients constitute the attacker cluster, and are blocked.
+Our algorithm depends on one single parameter $\delta$ which can be dynamically computed as a proportion of the size of the error space for the current round, such as:
+$\delta=\lambda(max(\epsilon_{u}^{t})-min(\epsilon_{u}^{t}))_{0<u<U}$ (3)
+Where $\lambda\in[0,1]$ is a hyper-parameter fixed in a data-agnostic manner at the start of the defense system. The full pseudo-code for the trust propagation algorithm is detailed in Alg. 1.
+Figure 4: Histogram of reconstruction error values measured for benign clients and attackers performing a Sign Flipping attack on MNIST, considering a trigger set $T_{s}$ of 250 samples.
+By sequentially building the trusted clients cluster rather than fixing a threshold, our detection handles attacks but also normal scenarios where the highest error clients are still benign.
+We observed that benign clients are very rarely misclassified with our method, excluding less than 1% of benign clients in the worst case (see 5.4).
 Algorithm 1 Trust propagation
-$\delta$, $clients\_sorted$, $sorted\_re$
-$benign\_clients\leftarrow clients\_sorted[0]$
-$i\leftarrow 0$
-$i_{max}\leftarrow\text{len}(clients\_sorted)-1$
-while ($i\neq i_{max}$) and ($sorted\_re[i+1]<sorted\_re[i]+\delta$do
-     $benign\_clients.\text{append}(clients\_sorted[i+1])$
-     $i\leftarrow i+1$
+[…]
+$\text{benign\_clients}\leftarrow\text{clients\_sorted[0]}$
+$\text{i}\leftarrow\text{0}$
+$\text{i\_{max}}\leftarrow\text{len(clients\_sorted)-1}$
+[…]
+  benign_clients.append(clients_sorted[i+1])
+  $\text{i}\leftarrow\text{i + 1}$
 end while
-return $benign\_clients$
+return  benign_clients
 5 Evaluation
@@
 5.1.1 FL setting
-We evaluate our proposition on Non-IID image classification tasks using two public datasets : MNIST [17] and FashionMNIST [51] datasets.
-The FL is orchestrated by a server that applies FedCLEAN as defense $D$ before each aggregation step with 100 total clients.
+We evaluated FL-CLEANER on Non-IID image classification tasks using two public datasets : MNIST [60] and FashionMNIST [61].
+In the following, FL is orchestrated by a server that applies FL-CLEANER as defense $D$ before each aggregation step with 100 total clients.
 50% of clients are randomly selected to participate in training at each round, and we use FedAVG [8] to aggregate model updates.
 As global model $M_{G}$, we use a standard CNN architecture that achieves good results on both datasets in the absence of attackers.
-In all our experiments, we set 30% of the client population to be attackers.
+In all our experiments, we set 30% of the client population to be attackers, in accordance with previous works in literature [16].
 5.1.2 Non-IID Data distribution
-We distribute the samples of MNIST and FashionMNIST datasets among clients to synthesize a Non-IID data distribution.
-As mentioned above, we craft distributions that achieve label skew and quantity skew according two scenarios: like in [31] and [54] based on a latent Dirichlet distributions; and following our own custom Non-IID distribution where clients are only allowed to hold samples from two classes.
-More in details, these two secarios are such as:
-  • Dirichlet distribution - we synthesize a population of $C$ clients by drawing a probability vector $\mathbf{q_{n}}\sim D_{C}(\alpha)$ for each class $n\in[1,N]$.
+Figure 5: Examples of our Non-IID distributions of the MNIST training samples in a scenario with 20 total clients. Colors represent the image classes present in the data.
+To synthesize a Non-IID data distribution among clients, we distribute the samples of MNIST and FashionMNIST datasets to achieve:
+  • Label skew based on a latent Dirichlet distributions as in [62] and [63] - we synthesize a population of $C$ clients by drawing a probability vector $\mathbf{q_{n}}\sim D_{C}(\alpha)$ for each class $n\in[1,N]$.
 Each client is allocated a portion $q_{n,c}$ of the total samples in class $n$.
 We set $\alpha=1$ [62].
-  • Our custom distribution - we enforce the Non-IID condition originally proposed in [41] where each client only possesses samples from two classes.
+  • Quantity skew based on a custom distribution we propose where clients are only allowed to hold samples from two classes - we enforce the Non-IID condition originally proposed in [8] where each client only possesses samples from two classes.
 This is achieved by sampling the full datasets using an inverse law with a scale factor of $\alpha$ and an offset of $\gamma$.
@@
 We set $\alpha=2000$, $\gamma=20$, and $r=2$.
-Fig. 3 shows an example of such distributions for the MNIST dataset.
-5.1.3 Defense setting
+Figure 5 shows an example of such distributions for the MNIST dataset.
+5.1.3 FL-CLEANER defense settings
 We use a CVAE conditioned on class labels.
@@
 The server $S$ trained its global model copy for 10 warmup epochs, then saved the activation maps from the next 10 epochs with a trigger set containing 250 samples extracted from the test set.
-We perform trust propagation with $\lambda=0.3$.
-Figure 3: Non-IID distributions of the MNIST training samples among 20 clients, colors represent target classes
+We perform trust propagation with $\lambda=0.3$ (see Section 4.3).
+In all our tests, we use all the layers of the model to extract activation maps.
 5.2 Considered Byzantine attacks
-Four attacks tested and parametrized in the literature were tested [12, 24, 6]:
-  • Sign-Flipping attacks: Attackers reverse the signs of their model updates by multiplying them by a factor of $-\xi$.
+Four attacks inspired from literature were tested [11, 12, 16]:
+  • Sign-Flipping attacks: Attackers reverse the signs of their model updates by multiplying them by a factor of $\xi$.
 This is akin to reversing gradient descent and leading the model in the direction of errors.
@@
   • Same Value attacks: malicious clients send a constant value $c$ as weight updates: $\omega^{\prime}_{i}=c$}.
-  • Scaling attack: malicious clients perform local SGD, then simply send back a scaled version of their updates, i.e., ${\omega^{\prime}_{i}=a\omega}$ with $a>1$
-Attacks’ parameters $\xi$, $\sigma$, $a$, and $c$ serve as amplification factors that increase the effect of the attack at the cost of making it more detectable For our tests to be nontrivial, we worked with non-upscaled attacks that send updates close to those of benign clients by fixing $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$.
-We will see that with no defense, these attacks are successful.
+  • Scaling attack: malicious clients perform local stochastic gradient descent, then simply send back a scaled version of their updates, i.e., $\omega^{\prime}_{i}=a\omega$} with $a>1$
+The above parameters $\xi$, $\sigma$, $a$, and $c$ control the force or the amplitude of the corresponding attack striking the compromise in between attack efficiency and detectability.
+For our tests to be nontrivial in terms of detection, we worked with attacks that send updates close to those of benign clients by fixing $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$.
+We will see that with no defense, these attacks are sufficient to collapse the global model, and that that they are also able to bypass detection by other defenses in literature (see Section 5.4).
+5.3 Considered backdoor attacks
+For our experiment, we choose two state-of-the-art FL backdoor attacks, which have proven to be successful in non-IID scenarios: The distributed backdoor attack (DBA) [64] and Neurotoxin [65].
+Let us recall that backdoor attacks work by adding a pattern (trigger) to training samples, changing their label to a target class, then training on these samples and sending back the corresponding model to the server.
+The two backdoors we tested work as follows :
+  • Distributed backdoor attacks (DBA) [64] : Malicious clients work together to inject the trigger, by each adding a part of the pattern to their samples.
+When the resulting models are aggregated by the server, the global model associates the entire trigger with the target class
+  • Neurotoxin [65] : When training on the poisoned samples, the attackers only update the smallest $k\%$ components of the gradient, which are less likely to be overwritten during aggregation.
+We test these attacks with a $10\times 10$ pixels white square pattern for both MNIST and FashionMNIST, split into four $5\times 5$ squares for DBA. For Neurotoxin, we set $k=95\%$.
 5.4 Experimental results
-Table 1: Defense results with the custom distribution
-MNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 93.69 84.92 92.78 92.19 91.91 FedCAM 93.02 92.09 93.97 11.29 92.83 FedCVAE 80.82 77.15 11.32 18.68 50.23 FedGuard 61.35 17.67 10.26 11.32 10.26 FLEDGE 93.30 88.06 92.56 92.04 14.19 NoDefense 93.61 18.34 11.32 11.32 12.00 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 50.60% 0.00% FedCVAE - 0.00% 50.00% 44.60% 1.42% FedGuard - 26.89% 22.15% 12.84% 17.46% FLEDGE - 0.00% 0.00% 0.00% 70.93% NoDefense - - - - - FedCLEAN 0.08% 0.00% 0.00% 0.00% 0.00% FedCAM 47.96% 0.00% 0.00% 47.12% 0.00% FedCVAE 46.79% 0.00% 43.43% 44.60% 1.67% FedGuard 82.01% 69.28% 75.33% 87.28% 82.51% FLEDGE 4.17% 0.00% 0.00% 0.00% 35.08% NoDefense - - - - - FshMNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 69.99 64.52 66.83 66.92 65.83 FedCAM 70.06 60.47 70.83 10.03 57.17 FedCVAE 65.02 64.93 10.07 17.92 60.30 FedGuard 17.29 20.08 10.04 10.04 10.04 FLEDGE 68.78 67.82 74.39 69.95 10.09 NoDefense 71.14 18.86 10.07 13.81 11.97 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 51.40% 0.00% FedCVAE - 0.00% 51.11% 52.05% 2.68% FedGuard - 27.43% 11.86% 24.11% 18.81% FLEDGE - 0.00% 0.00% 0.00% 70.36% NoDefense - - - - - FedCLEAN 0.68% 0.00% 0.00% 0.00% 0.00% FedCAM 47.48% 0.00% 0.00% 49.51% 0.00% FedCVAE 45.61% 0.00% 43.01% 45.20% 5.98% FedGuard 95.91% 70.90% 83.49% 73.33% 80.52% FLEDGE 3.38% 0.00% 0.00% 0.00% 27.78% NoDefense - - - - -
-Table 2: Defense results with the Dirichlet distribution
-MNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 95.52 95.39 95.39 95.42 95.51 FedCAM 95.71 95.59 95.57 10.29 95.44 FedCVAE 94.56 93.86 11.32 10.12 17.05 FedGuard 93.20 16.69 11.32 10.33 13.93 FLEDGE 95.58 95.61 95.82 95.31 8.59 NoDefense 95.76 15.63 11.32 10.33 9.76 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 53.15% 0.00% FedCVAE - 0.00% 56.64% 60.43% 6.54% FedGuard - 21.37% 43.48% 34.14% 45.23% FLEDGE - 0.00% 0.00% 0.00% 50.81% NoDefense - - - - - FedCLEAN 0.04% 0.00% 0.00% 0.00% 0.00% FedCAM 48.04% 0.00% 0.00% 47.97% 0.00% FedCVAE 42.82% 0.00% 49.30% 46.54% 4.90% FedGuard 55.99% 77.73% 58.64% 66.06% 53.67% FLEDGE 0.04% 0.00% 0.00% 0.00% 39.07% NoDefense - - - - - FshMNIST NAtk AdNse SmVal SgnFlp Scaled FedCLEAN 70.86 70.47 68.99 71.06 71.85 FedCAM 74.37 72.55 71.43 10.03 70.19 FedCVAE 68.08 68.09 10.07 17.89 66.75 FedGuard 60.03 15.92 10.07 10.07 12.13 FLEDGE 71.75 70.71 70.40 70.55 9.06 NoDefense 74.39 17.88 10.07 10.01 11.95 FedCLEAN - 0.00% 0.00% 0.00% 0.00% FedCAM - 0.00% 0.00% 52.39% 0.00% FedCVAE - 0.00% 49.35% 56.55% 2.08% FedGuard - 34.65% 31.53% 29.27% 30.50% FLEDGE - 0.00% 0.00% 0.00% 85.71% NoDefense - - - - - FedCLEAN 0.12% 0.00% 0.00% 0.00% 0.00% FedCAM 48.72% 0.00% 0.00% 48.91% 0.00% FedCVAE 44.29% 1.08% 41.91% 45.92% 4.21% FedGuard 70.00% 66.30% 70.67% 72.55% 70.22% FLEDGE 2.59% 0.00% 0.00% 0.00% 13.44% NoDefense - - - - -
-Tables 1 and 2 report the results of testing on different distributions of the MNIST and FashionMNIST datasets.
-We use three standard metrics to measure the effectiveness of each defense system: the global model’s accuracy on the test dataset (ACC), the attacker detection false negative (FN), which refers to the fraction of attackers that manage to get through the defense system, and benign detection false positive (FP) which represents the fraction of blocked benign clients.
-For a byzantine attack, the defense requirement R1 requires FN to be 0% in order for the defense to be considered successful.
-The defense also needs to not disrupt the training (R2), which means ACC should be as close as possible to the benchmark no defense and no attack accuracy, and FP should be as close as possible to 0%.
-As we can observe, FedCLEAN is the only defense system that achieves 0% of FN against all types of Byzantine attacks in all tested scenarios.
-We also see that when attackers are present, FedCLEAN does not block any benign clients (i.e FP=0%).
+Figure 6: Defense Results on MNIST considering two non-IID distributions against Byzantine attacks: Additive noise, Same value, Sign Flipping and Scaled attacks of parameters $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$ respectively.
+Figure 7: Defense Results on FashionMNIST considering two non-IID distributions against Byzantine attacks: Additive noise, Same value, Sign Flipping and Scaled attacks of parameters $\xi=1$, $\sigma=0.1$, $c=0.01$ and $a=10$ respectively.
+For Byzantine attacks, we use three standard metrics to measure the effectiveness of each defense system: the global model’s accuracy on the test dataset (ACC); the attacker detection recall (Recall) which refers to the percentage of attackers successfully blocked by the defense system; and, benign detection false positive (FPR) which represents the fraction of blocked benign clients.
+Figures 6 and 7 report the results of testing on the two distributions of the MNIST and FashionMNIST datasets.
+For byzantine attacks, recall that a single attacker is sufficient to collapse the global model in just one FL round.
+Thus, a Recall value of 100% is required in order the defense to be considered successful and reach the defense requirement R1 (see Section 3.2).
+The defense also needs to not disrupt the training (R2), which means ACC should be as close as possible to the benchmark no defense and no attack accuracy, and FPR should be as close as possible to 0%.
+As we can observe in Figures 6 and 7, compared to the most recent effective methods in literature (see Section 1.1), FL-CLEANER is the only defense system that achieves 100% Recall against all types of Byzantine attacks in all tested scenarios.
+We can also see that when attackers are present, FL-CLEANER does not block any benign clients (i.e FPR=0%).
 When attackers are absent, it misclassifies less than 1% of benign clients.
-Additionally, across 20 tested scenarios, FedCLEAN achieves the highest test accuracy in 10 scenarios and matches benchmark accuracy in all 20.
-But one should note that in all 10 scenarios where FedCLEAN does not achieve the highest accuracy, it achieves 100% attacker detection and 0% or close misclassification, which suggests that the gap in accuracy is due to the random attribution of samples at the start of each experiment.
+Additionally, across 20 tested scenarios, FL-CLEANER achieves the highest test accuracy in 10 scenarios and matches benchmark accuracy in all 20.
+Note that one reason why FL-CLEANER does not reach the highest accuracy achieved in the scenario with no-defense and no attack may come from the fact that in the experiments of FL-CLEANER only 70% of the training data are used for training.
+The rest is used to simulate malicious clients for test purposes while in the scenario with no-defense and no attack, all data are used for training.
+Regardless, FL-CLEANER achieves 100% attacker detection and 0% or close miss-classification in each of these cases.
+It is important to note that other defenses were implemented according to the parameters provided by their authors.
+For FedCVAE for example, this meant using the whole test set, i.e, 10 000 images, to train the its CVAE.
+It can be seen that FL-CLEANER achieves better performance with a much smaller trigger set of 250 samples.
+Defense DBA Neurotoxin
+ASR ($\downarrow$) ACC ($\uparrow$) ASR ($\downarrow$) ACC ($\uparrow$)
+NoDefense 95.4% 21.2% 9.1% 92.8%
+FLAME 0.0% 9.8% 0.0% 16.8%
+FLDetector 11.9% 85.7% 9.6% 93.9%
+FL-CLEANER 0.2% 95.0% 0.0% 95.1%
+Table 1: Attack Success Rate (ASR) and Accuracy (ACC) for different defenses and attacks on MNIST. Arrows represent ideal outcome.
+Defense DBA Neurotoxin
+ASR ($\downarrow$) ACC ($\uparrow$) ASR ($\downarrow$) ACC ($\uparrow$)
+NoDefense 6.1% 65.8% 100% 59.6%
+FLAME 9.1% 42.1% 23.4% 54.1%
+FLDetector 16.3% 75.0% 43.9% 71.4%
+FL-CLEANER 0.2% 67.2% 0.3% 69.1%
+Table 2: Attack Success Rate (ASR) and Accuracy (ACC) for different defenses and attacks on FashionMNIST. Arrows represent ideal outcome.
+Concerning backdoor attacks, let us recall that the goal of the attacker is to achieve high ASR while preserving the model’s performance on the main task (see Section 3.1).
+Thus, we evaluate our defense on the impact it has on this metric.
+To evaluate this, we first test the backdoors in the absence of our defense, and track their ASR over the training epochs.
+Then, we do the same with a defense applied to compare.
+In addition, we report the final attacker success rate and test accuracy at the end of the run in separate tables.
+In our testing, attackers struggled to insert the backdoor when training on the more restrictive custom distribution.
+So we focused our tests on the Dirichlet distribution.
+For comparison with other defenses, we decided not to take into account backdoor defenses the strategy of which includes a tolerance mechanism sometimes letting one to several attackers get through the defense [27, 15].
+These approaches are by design not byzantine-robust.
+Indeed, most byzantine attacks only need one attacker to get through.
+We choose to focus on two backdoor defenses that filter out all detected attackers : FLAME [28] and FLDetector [29].
+They are also considered as the state-of-the-art [66]
+Figures 9 and 9 show the evolution of backdoor accuracy (ASR) over a 50-round FL run.
+Tables 1 and 2 report the final attacker success rate (ASR) and test accuracy (ACC) observed at the end of these runs.
+We see that :
+  1. A
+    FL-CLEANER cancels the effect of both attacks : After a few rounds, we see that it becomes impossible for the attacker to maintain an ASR not nearing 0%.
+This validates the backdoor defense requirement R3.
+  2. B
+    FL-CLEANER is the only effective defense that does not harm the model accuracy : As seen in Tables 1 and 2, when other defenses manage to mitigate the backdoor attack, it comes at the cost of the model’s accuracy.
+This is due to the defenses blocking too many benign clients.
+Which makes FL-CLEANER the only defense to validate the defense requirement R2.
+Figure 8: Backdoor Attacker Success rates on MNIST
+with and without defense (smoothed)
+Figure 9: Backdoor Attacker Success rates on Fashion-
+MNIST with and without defense (smoothed)
 6 Conclusion & future work
-We introduced FedCLEAN, an autoencoder-based defense system designed to identify malicious clients in Non-IID federated learning environments from updated models’ activation maps.
-With two core components: A framework for training CVAEs for an FL defense system leveraging model activation maps normalized with GeoMed, DKL annealing, and a realistic trigger set requirement.
-And a dynamic one-parameter client selection algorithm that minimizes the exclusion of benign clients.
-We demonstrate that, in image classification benchmarks, FedCLEAN is robust against Byzantine attacks under a Non-IID client data distribution and achieves top performance in benign client missclassification tests.
-Future work will focus on broader machine learning applications, backdoor attacks, as well as improving our method by enhancing CVAE training through adaptive regularization, and implementing a fully non-parametric trust propagation algorithm.
-6.0.1 Acknowledgements
+In this paper, we introduced FL-CLEANER the first method robust against Byzantine and backdoor attacks under a Non-IID client data distribution.
+Its originality stands on an autoencoder-based defense system designed to identify malicious clients in Non-IID federated learning environments from updated models’ activation maps with other core components: a framework for training CVAEs for an FL defense system leveraging model activation maps normalized with GeoMed, DKL annealing, and a realistic trigger set requirement.
+Another strength of our proposal is that it depends on a dynamic one-parameter client selection algorithm that minimizes the exclusion of benign clients.
+We demonstrated that, in image classification benchmarks, FL-CLEANER is robust against Byzantine attacks as well as effective to mitigate tested state-of-the-art backdoor attacks, achieving top performance in benign client misclassification tests compared to current defenses from the literature.
+Future work will focus on broader machine learning applications, as well as improving our method by enhancing CVAE regularization in the training, and implementing an improved trust propagation algorithm.
+Acknowledgments
 This work was partly supported by the Inserm industrial chair CYBAILE and the PEPR digital health SSF-ML-DH project, ANR-13- LAB2-0007-01